. Total Wine is planning to open a megastore in a suburban town. There are currently 6 small stores selling 1.5L bottles of wine. Each store experiences identical supply/demand behavior: Mean demand of 54 bottles/week. The standard deviation of the demand 13 bottles/week is given to you The supply lead time is 4.2 weeks, and the target service level is 90%. () What is the safety stock for each of the small stores? Answer Check i. What is the Big-Box safety stock advantage of Total Wine? Answer: Check
Answer #1

Answer i:

  • 43.69299

Explanation:

  • Z value for 90% CI = 1.64
  • Weekly demand = D = 54
  • Standard deviation per store =S.D. = 13 bottles/week
  • Lead time = LT = 4.2 weeks
  • Safety stock for each = Z * S.D. * (LT)1/2 = 1.64*13*(4.2)1/2= 43.69299

Answer ii:

  • 155.1839

Explanation:

  • Z value for 99% CI = 1.64
  • Weekly demand of big box = D*6 = 324
  • Standard deviation for big box =S.D. * (6)^0.5 = 31.843366bottles/week
  • Lead time = LT = 4.2 weeks
  • Safety stock for each = Z * S.D. * (LT)1/2 = 1.64*13*(4.2)1/2= 107.0255
  • Total safety stock in 6 small stores = 43.6999*6 = 262.1994
  • Advantage of big box = 262.1994 -107.0255 = 155.1739